Evaluate the integral of dx/(x ln x)


1
Expert's answer
2021-03-29T13:44:25-0400
"\int \dfrac{dx}{x \ln x}"

Use "u"-substitution


"u=\ln x, du=\dfrac{1}{x}dx"

"\int \dfrac{dx}{x \ln x}=\int \dfrac{du}{u}=\ln|u|+C=\ln{(|\ln x|)}+C"


"\int \dfrac{dx}{x \ln x}=\ln{(|\ln x|)}+C"